# Additivity of Crossing Number via Restricted Reidemeister Moves

### v1

@@ 6.2 Minimal Splitting Curves
Use the same argument as in the proof of Claim 6.17.3 to show that $A^{\prime}\setminus U\subset A\setminus U$. Suppose $z\in A^{\prime}\cap U$. Then $h(z)$ cannot be between the lines $\alpha^{+}$ and $\alpha^{-}$ (Figure 6) because then we would be able to find a point in hAhA which contradicts Claim 6.17.2. So $h(z)\in[0,1]^{2}$ or $h(z)\in[-1,0]^{2}$. Suppose $h(z)\in[0,1]^{2}$ and let $\gamma$ be a path connecting $z$ to $H^{+}$. By the same argument as above, $h\left[\gamma\cap U\right]$ cannot pass between $\alpha^{+}$ and $\alpha^{-}$ and also $\gamma\cap S^{\prime}=\varnothing$. But this implies that $\gamma$ in fact witnesses that $z\in A$. The case $h(z)\in[-1,0]^{2}$ follows in the same way. ∎
Recall that $J_{0}$ was defined to witness that $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$ and lies outside $U$, so also witnesses $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\leqslant p+1$. Thus, it is enough to show that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\geqslant p+1$. Suppose $J$ is a vertical splitting curve such that $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})$. If $J$ does not intersect $U$, then it is a splitting curve for $S$, so $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{\mathbf{i}}(S,J)\geqslant p+1$. Otherwise, there is $t$ with $J(t)\in U$. Let $t$ be the maximum such $t$. W.l.o.g. by a general position argument, $J(t)\notin|S^{\prime}|$. If $J(t)\in A^{\prime}$, let $\gamma$ be a curve connecting $H^{+}$ to $J(t)$ which lies in $A^{\prime}$. By Claim 6.17.4 it lies in AA, and so outside $S$. Let $J^{\prime}=\gamma{}^{\frown}(J\!\restriction\![t,1])$.
@@
For any vertical splitting curve $J$ we have $\operatorname{\mathbf{i}}(hT,J)=\operatorname{\mathbf{i}}(T,h^{-1}J).$ If $J$ is a vertical splitting curve, then so is $h^{-1}J$, by (6.34). Thus, $\operatorname{MIN}_{\operatorname{vert}}(hT)\geqslant\operatorname{MIN}_{\operatorname{vert}}(T)=k-1$ and hThT is of type $(0,k,0,k)$. By Lemmas 6.18 and 6.8 we have $\operatorname{MIN}_{\operatorname{vert}}(hT)=k$. Let $\beta$ be the $k$-bridge given by Lemma 6.17. Then $\Sigma_{1}=h^{-1}\beta$ is as desired. ∎
@@
Let $J$ be an ijij-splitting curve for $T$. W.l.o.g. $i\leqslant j$. If $i=j$, $w(i,j)=0$, so the statement is trivial. So we are left with the case $i<j$. Suppose first that $i<j\leqslant k$. Since $T$ is of type (1,k1,1,k1)(1,k-1,1,k-1), $L_{i}(T)$ and $L_{j}(T)$ are now subsets of $V^{-}$. We have to show that $\operatorname{\mathbf{i}}(T,J)\geqslant w(i,j)=|j-i|=j-i$. Suppose not. Let $J_{1}$ be the line from the top endpoint of $V^{-}$, $(0,1)$, to $J(0)$. Then $\operatorname{\mathbf{i}}(T,J_{1})=i-1$. Let $J_{2}$ be the line from $J(1)$ to the lower endpoint of $V^{-}$, $(0,0)$. Then $\operatorname{\mathbf{i}}(T,J_{2})=k-j$. Let $J_{*}=J_{1}{}^{\frown}J{}^{\frown}J_{2}$. Then $\operatorname{\mathbf{i}}(T,J_{*})=\operatorname{\mathbf{i}}(T,J_{1})+\operatorname{\mathbf{i}}(T,J)+\operatorname{\mathbf{i}}(T,J_{2})=i-1+\operatorname{\mathbf{i}}(T,J)+k-j<i-1+j-i+k-j=k-1.$ By pushing $J_{*}$ slightly away from $V^{-}$ into $\square$, obtain a vertical splitting curve $J_{*}^{\prime}$ with $\operatorname{\mathbf{i}}(T,J_{*}^{\prime})<k-1$, a contradiction. Same argument for $i>j>k$ noting that then $L_{i}(T),L_{j}(T)\subset V^{+}$. Suppose
@@ 7.1 Subdivided Tangle Diagrams
A subdivided t-diagram $(T,T^{+},T^{-},\{1\},M,\ell)$ is centered if the knot diagram defined by $T\!\restriction\!(I\times\{1\})$ is centered. The lower and upper minimal splitting numbers $\operatorname{MIN}_{\pm}(T^{\mp})$ are defined in the same way as for (partial) knot diagrams. Consistent with the terminology for knot diagrams, call $T^{\pm}$ the partial tangle diagrams, or partial t-diagrams. The $\sim^{*}_{\mathcal{R}}$, $\sim^{*}_{0}$, and $\sim_{\mathcal{R}}$-equivalences are extended to subdivided t-diagrams in an obvious way. A subdivided t-diagram is permitted if for all $x\in T^{+}\cap T^{-}$, $\ell^{-1}(x)\in(I^{+}\cup(S^{1}\times M_{+}))$. Extend the definitions of hThT and $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ (see Definition 4.4 and Section 4.4) for partial and subdivided tangle diagrams in an obvious way, but note that we will treat $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ as a t-diagram, not as a subdivided one, because it might not be of type $(1,0,1,0)$ as condition ($PT_{1}$) above requires. Also, if $T^{+}$ is a partial t-diagram, then $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ is a t-diagram when $o\notin G$. All situations where we will use this notion will be “non-problematic” in this sense.
@@ 7.3 Straightening Rays
The set $A=\square(2\delta)\setminus(\square(\delta)\cup J)$ is connected and $b_{-}(\delta)$ is on the boundary of AA. So there is $\gamma_{1}\subset A\cup\{b_{-}(\delta)\}$ which starts at $b_{-}(\delta)$ and ends at $J^{\prime}(0)$. The set $B=U(2\delta)\setminus(U(\delta)\cup J)$ is also connected and $d_{-}(\delta)$ is on the boundary of $B$, so there is a curve $\gamma_{2}\subset B\cup\{d_{-}(\delta\}$ which starts at $J^{\prime}(1)$ and ends at $d_{-}(\delta)$. Let $J_{1}=J\!\restriction\![0,1-\delta]$ and let
@@ 7.4 The Fan Lemma
Figure 11: Illustrating the maps $\psi$ and $\eta^{-1}\circ\psi\circ\eta$ from the proof of Lemma $7.$ Figure (a) is the domain and (b) the range of $\psi$. The colored regions are mapped respectively to the same colored regions. The light-red area is where $\psi$ is the identity. The yellow and orange areas are squeezed and the blue and green areas are stretched. The gray areas is where $\psi$ is defined by interpolating between the light-red and other colored areas. In (c) and (d) the domain and range of $\eta^{-1}\circ\psi\circ\eta$ is shown. The green and blue areas form an area inside of $\tau(\delta)$ according to (7.1). Figures (c) also shows the line segments $K^{+}\cap U(\delta)$ and $K^{+}\cap\square_{\delta}$ (red) as well as $K^{-}\cap U(\delta)$ and $K^{-}\cap\square_{\delta}$ (blue). In (d) the images of these line segments are shown. Notice how the gray area “drags” them into the lower region due to the interpolation.
@@
Let $\operatorname{Fan}(T)$ be the subdivided tangle diagram given by Lemma $7.$ Since $\operatorname{Fan}(T)\sim^{*}_{0}T$ and $\operatorname{MIN}_{\operatorname{ray}}$ is invariant under $\sim^{*}_{0}$ (Lemma 5.2 applies to composite tangle as well as knot diagrams), we can assume without loss of generality that $\operatorname{Fan}(T)=T$. Let $S^{-}$ and $S^{+}$ be as defined in the statement of Lemma $7.$ By (F4F{4}), (F6F{6}), and (F8F{8}), $S^{-}$ and $S^{+}$ satisfy the assumptions of Proposition 6.1, so it follows that

### v2

@@ 6.2 Minimal Splitting Curves
Use the same argument as in the proof of Claim 6.17.3 to show that $A^{\prime}\setminus U\subset A\setminus U$. Suppose $z\in A^{\prime}\cap U$. Then $h(z)$ cannot be between the lines $\alpha^{+}$ and $\alpha^{-}$ (Figure 6) because then we would be able to find a point in $hA$ which contradicts Claim 6.17.2. So $h(z)\in[0,1]^{2}$ or $h(z)\in[-1,0]^{2}$. Suppose $h(z)\in[0,1]^{2}$ and let $\gamma$ be a path connecting $z$ to $H^{+}$. By the same argument as above, $h\left[\gamma\cap U\right]$ cannot pass between $\alpha^{+}$ and $\alpha^{-}$ and also $\gamma\cap S^{\prime}=\varnothing$. But this implies that $\gamma$ in fact witnesses that $z\in A$. The case $h(z)\in[-1,0]^{2}$ follows in the same way. ∎
Recall that $J_{0}$ was defined to witness that $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$ and lies outside $U$, so also witnesses $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\leqslant p+1$. Thus, it is enough to show that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\geqslant p+1$. Suppose $J$ is a vertical splitting curve such that $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})$. If $J$ does not intersect $U$, then it is a splitting curve for $S$, so $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{\mathbf{i}}(S,J)\geqslant p+1$. Otherwise, there is $t$ with $J(t)\in U$. Let $t$ be the maximum such $t$. W.l.o.g. by a general position argument, $J(t)\notin|S^{\prime}|$. If $J(t)\in A^{\prime}$, let $\gamma$ be a curve connecting $H^{+}$ to $J(t)$ which lies in $A^{\prime}$. By Claim 6.17.4 it lies in $A$, and so outside $S$. Let $J^{\prime}=\gamma{}^{\frown}(J\!\restriction\![t,1])$.
@@
For any vertical splitting curve $J$ we have $\operatorname{\mathbf{i}}(hT,J)=\operatorname{\mathbf{i}}(T,h^{-1}J).$ If $J$ is a vertical splitting curve, then so is $h^{-1}J$, by (6.34). Thus, $\operatorname{MIN}_{\operatorname{vert}}(hT)\geqslant\operatorname{MIN}_{\operatorname{vert}}(T)=k-1$ and $hT$ is of type $(0,k,0,k)$. By Lemmas 6.18 and 6.8 we have $\operatorname{MIN}_{\operatorname{vert}}(hT)=k$. Let $\beta$ be the $k$-bridge given by Lemma 6.17. Then $\Sigma_{1}=h^{-1}\beta$ is as desired. ∎
@@
Let $J$ be an $ij$-splitting curve for $T$. W.l.o.g. $i\leqslant j$. If $i=j$, $w(i,j)=0$, so the statement is trivial. So we are left with the case $i<j$. Suppose first that $i<j\leqslant k$. Since $T$ is of type (1,k1,1,k1)(1,k-1,1,k-1), $L_{i}(T)$ and $L_{j}(T)$ are now subsets of $V^{-}$. We have to show that $\operatorname{\mathbf{i}}(T,J)\geqslant w(i,j)=|j-i|=j-i$. Suppose not. Let $J_{1}$ be the line from the top endpoint of $V^{-}$, $(0,1)$, to $J(0)$. Then $\operatorname{\mathbf{i}}(T,J_{1})=i-1$. Let $J_{2}$ be the line from $J(1)$ to the lower endpoint of $V^{-}$, $(0,0)$. Then $\operatorname{\mathbf{i}}(T,J_{2})=k-j$. Let $J_{*}=J_{1}{}^{\frown}J{}^{\frown}J_{2}$. Then $\operatorname{\mathbf{i}}(T,J_{*})=\operatorname{\mathbf{i}}(T,J_{1})+\operatorname{\mathbf{i}}(T,J)+\operatorname{\mathbf{i}}(T,J_{2})=i-1+\operatorname{\mathbf{i}}(T,J)+k-j<i-1+j-i+k-j=k-1.$ By pushing $J_{*}$ slightly away from $V^{-}$ into $\square$, obtain a vertical splitting curve $J_{*}^{\prime}$ with $\operatorname{\mathbf{i}}(T,J_{*}^{\prime})<k-1$, a contradiction. Same argument for $i>j>k$ noting that then $L_{i}(T),L_{j}(T)\subset V^{+}$. Suppose
@@ 7.1 Subdivided Tangle Diagrams
A subdivided t-diagram $(T,T^{+},T^{-},\{1\},M,\ell)$ is centered if the knot diagram defined by $T\!\restriction\!(I\times\{1\})$ is centered. The lower and upper minimal splitting numbers $\operatorname{MIN}_{\pm}(T^{\mp})$ are defined in the same way as for (partial) knot diagrams. Consistent with the terminology for knot diagrams, call $T^{\pm}$ the partial tangle diagrams, or partial t-diagrams. The $\sim^{*}_{\mathcal{R}}$, $\sim^{*}_{0}$, and $\sim_{\mathcal{R}}$-equivalences are extended to subdivided t-diagrams in an obvious way. A subdivided t-diagram is permitted if for all $x\in T^{+}\cap T^{-}$, $\ell^{-1}(x)\in(I^{+}\cup(S^{1}\times M_{+}))$. Extend the definitions of $hT$ and $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ (see Definition 4.4 and Section 4.4) for partial and subdivided tangle diagrams in an obvious way, but note that we will treat $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ as a t-diagram, not as a subdivided one, because it might not be of type $(1,0,1,0)$ as condition ($PT_{1}$) above requires. Also, if $T^{+}$ is a partial t-diagram, then $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ is a t-diagram when $o\notin G$. All situations where we will use this notion will be “non-problematic” in this sense.
@@ 7.3 Straightening Rays
The set $A=\square(2\delta)\setminus(\square(\delta)\cup J)$ is connected and $b_{-}(\delta)$ is on the boundary of $A$. So there is $\gamma_{1}\subset A\cup\{b_{-}(\delta)\}$ which starts at $b_{-}(\delta)$ and ends at $J^{\prime}(0)$. The set $B=U(2\delta)\setminus(U(\delta)\cup J)$ is also connected and $d_{-}(\delta)$ is on the boundary of $B$, so there is a curve $\gamma_{2}\subset B\cup\{d_{-}(\delta\}$ which starts at $J^{\prime}(1)$ and ends at $d_{-}(\delta)$. Let $J_{1}=J\!\restriction\![0,1-\delta]$ and let
@@ 7.4 The Fan Lemma
Figure 11: Illustrating the maps $\psi$ and $\eta^{-1}\circ\psi\circ\eta$ from the proof of Lemma 7.7. Figure (a) is the domain and (b) the range of $\psi$. The colored regions are mapped respectively to the same colored regions. The light-red area is where $\psi$ is the identity. The yellow and orange areas are squeezed and the blue and green areas are stretched. The gray areas is where $\psi$ is defined by interpolating between the light-red and other colored areas. In (c) and (d) the domain and range of $\eta^{-1}\circ\psi\circ\eta$ is shown. The green and blue areas form an area inside of $\tau(\delta)$ according to (7.1). Figures (c) also shows the line segments $K^{+}\cap U(\delta)$ and $K^{+}\cap\square_{\delta}$ (red) as well as $K^{-}\cap U(\delta)$ and $K^{-}\cap\square_{\delta}$ (blue). In (d) the images of these line segments are shown. Notice how the gray area “drags” them into the lower region due to the interpolation.
@@
Let $\operatorname{Fan}(T)$ be the subdivided tangle diagram given by Lemma 7.7. Since $\operatorname{Fan}(T)\sim^{*}_{0}T$ and $\operatorname{MIN}_{\operatorname{ray}}$ is invariant under $\sim^{*}_{0}$ (Lemma 5.2 applies to composite tangle as well as knot diagrams), we can assume without loss of generality that $\operatorname{Fan}(T)=T$. Let $S^{-}$ and $S^{+}$ be as defined in the statement of Lemma 7.7. By (F4F{4}), (F6F{6}), and (F8F{8}), $S^{-}$ and $S^{+}$ satisfy the assumptions of Proposition 6.1, so it follows that